OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (x - 2)(x - 1)(x - 6). 2/x-2

jimthompson5910 (jim_thompson5910):

\[\frac{2}{x-2}\] \[\frac{2(x-1)(x-6)}{(x-2)(x-1)(x-6)}\] \[\frac{2(x^2-7x+6)}{(x-2)(x-1)(x-6)}\] \[\frac{2x^2-14x+12}{(x-2)(x-1)(x-6)}\] So \[\frac{2}{x-2}=\frac{2x^2-14x+12}{(x-2)(x-1)(x-6)}\]